# Topological columnar nano-SQUID based on a 3D topological insulator

### v2

@@ Abstract
Finding an ideal platform to generate non-Abelian Majorana zero-modes (MZMs) is a crucial first step for topological quantum computing. A 3D topological insulator (TI) is promising for this purpose due to the spin-momentum-locked surface state. Here we propose a novel yet simple TI platform that gives rise to a robust topological phase and can realize a well-defined fermion parity necessary for quantum computing. It consists of a bulk-insulating rectangular TI nanowire laterally sandwiched by two superconductors. In this structure, the top and bottom surfaces individually work as SNS line junctions, forming a nanometer-scale columnar SQUID in which the nanowire cross-section defines the threading flux $\Phi$ in axial magnetic fields. We demonstrate that a TI device of this structure indeed presents SQUID-type oscillations of the critical current $I_{c}$ as a function of $\Phi$ with period $\Phi_{0}^{s}=\frac{h}{2e}$. It can be tuned by a back gate to attain a full $I_{c}$ suppression and then a nearly-full recovery as a function of $\Phi$, an indication that the superconductivity is completely surface-dominated. Our theory shows that, when the two junctions are asymmetric, a robust topological phase hosting MZMs occurs periodically for $(n-\frac{1}{2})\Phi_{0}^{s}<\Phi<(n+\frac{1}{2})\Phi_{0}^{s}$ with odd-integer $n$.
@@ I Introduction
The main feature of three-dimensional (3D) topological insulators (TIs) is the existence of gapless surface Dirac fermions that are protected by time-reversal symmetry [1, 2, 3]. These Dirac fermions are spin-non-degenerate with the spin direction dictated by the momentum; this spin-momentum locking is the key ingredient for realizing “spinless” topological superconductivity that gives rise to Majorana zero-modes (MZMs) [4]. The non-Abelian statistics obeyed by MZMs allows for topological quantum computing based on braiding operations [5], so TIs have a high potential for such applications. However, TIs have been left behind in the main-stream efforts to realize topological quantum computing, largely due to the less developed technologies for TI device fabrications compared to semiconductor devices [6]. Unfortunately, despite significant efforts in the main-stream semiconductor platforms, the generation of MZMs is still elusive [7, 8]. This is mainly due to the sensitivity of the semiconductor platforms to Coulomb disorder [9, 8, 10, 11], and a new platform that is inherently insensitive to chemical-potential fluctuations is strongly called for.
@@
Such a topological 1D Majorana state should in principle have a MZM when it has an end, but if the line junction ends in the middle of the TI surface, the MZM will hybridize with the gapless Dirac fermions in the rest of the surface; therefore, Fu and Kane proposed a phase-biased tri-junction to generate a MZM that is protected by a gap [4]. Even though a protected MZM can be generated in this way, when the tri-junction is made on the top surface of a TI and the bottom surface is left normal (which is expected when the TI thickness is more than $\sim$20 nm [12, 13]), the gapless Dirac fermions at the bottom surface may compromise the parity of the device, causing problems for a Majorana qubit that encodes quantum information through fermion parity [14]. Similar problem exists when using the MZM appearing at the Josephson vortex core of a TI line junction [15, 16]. Therefore, for topological quantum computing [5], gapping out the whole surface is important [17]. If this is achieved, the MZMs generated in a junction by phase biasing is robust and disorder-resilient [9], as their existence does not depend on the chemical potential at all provided that the TI is bulk-insulating.
@@
Figure 1: Sandwich TINW Josephson junction. (a) False-color SEM image of device A; an exfoliated TI flake is dry-etched to leave a NW (pink), and the etched area is subsequently filled with Nb (green), such that the TINW is side-contacted by superconducting Nb. There are remaining TI pieces that do not play any role. (b) Schematic of our device: Back-gated TINW is laterally sandwiched between two superconducting Nb electrodes to a Josephson junction. In- and out-of-plane magnetic fields can be applied. (c) $dV/dI$ vs $I_{\rm b}$ data in device A at various $T$ shown for zero back-gate voltage; inset shows the $T$-dependence of the critical current $I_{c}$ extracted from these data. We define $I_{c}$ as the bias current above which $dI/dV$ exceeded 50 $\Omega$. (d) Differential resistance $dV/dI$ vs junction voltage $V$ for device A in $B$ = 0 T. Red arrows mark the position of the peaks due to multiple Andreev reflection (MAR). We assign the index $n$ from 1 to 4 to the peaks ($n$ = 1 for the largest $V$). Inset shows a plot of the $V$ value for index $n$ (which we call $V_{n}$) vs $1/n$. (e-g) Color mappings of $dV/dI$ as a function of $I_{\mathrm{b}}$ and $B_{\mathrm{z}}$ for devices A, B and C measured without any in-plane magnetic field at $T=30$ mK, presenting Fraunhofer patterns. The critical current manifests itself as a bright orange line to border the black area corresponding to zero resistance. Irregular Fraunhofer features are attributed to an inhomogenous supercurrent density along the junction.
@@
To circumvent the problems and exploit the benefits of the TI line junction, we fabricate unconventional devices in which a rectangular nanowire, etch-fabricated from a bulk-insulating TI flake [18], is side-contacted by a superconductor on both left and right sides, creating a lateral sandwich SNS junction with a TI nanowire (TINW) as the N part (see Fig. 1b). In such a sandwich junction, if the TINW is bulk-insulating, the supercurrent flows only through the top and bottom surfaces (and a very small portion through the ends of the TINW) via the ABSs formed there. The whole surface of this TINW is gapped, with the proximity-induced SC gap in the side surfaces and the mini-gap of the ABSs elsewhere [4]. Since the axial magnetic field applied parallel to the TINW creates a magnetic flux $\Phi$ which controls the phase difference $\varphi$ at the top and bottom JJs, this structure can be viewed as a nanometer-scale columnar SQUID (superconducting quantum interference device) where the loop area corresponds to the TINW cross-section. A similar nano-SQUID was previously realized with a double-layer graphene heterostructure [19]. Our device has a bottom gate, which allows us to control the chemical potential in the TINW as well as to introduce asymmetry between the top and bottom junctions. A similar lateral sandwich junction was previously realized with a TI by using a diffusion process [20], but the device was not gate-tunable and it was not clear if the supercurrent flows only through the surface or also through the bulk.
@@
In our devices, we find that the critical current $I_{c}$ of the device exhibits pronounced oscillations as a function of $\Phi$ with a period $h/(2e)$ and, in certain situations, $I_{c}$ is almost completely suppressed when $\Phi=h/(4e)$ but recovers almost entirely for $\Phi=h/(2e)$. This is exactly the behavior expected for a nano-SQUID with completely surface-dominated supercurrent with negligible bulk contribution. Theoretically, we first discuss a phenomenological SQUID model to explain the observed behavior, and then present numerical simulations based on a microscopic model of two line junctions made on a TINW surface, which support the SQUID model. Furthermore, we predict that when the SQUID is asymmetric, the magnetic flux $\Phi$ causes only one of the two line junctions to go through the topological phase transition of the Fu-Kane type [4] and the SQUID is topological for $(n-\frac{1}{2})\Phi_{0}^{s}<\Phi<(n+\frac{1}{2})\Phi_{0}^{s}$ with odd-integer $n$. Note that there have also been related theoretical proposals to use the quasi-one-dimensional states of TINWs under superconducting proximity effect for generating MZMs [21, 22, 23, 24, 25]; however, these rely on the realization of a well-defined subband structure, i.e., that electron states on the top and bottom surfaces are coherently connected, which is most likely not realized in our device.
@@
II Results
Device characterization: Our TINWs are based on bulk-insulating BiSbTeSe2 [26, 27] and are side-contacted by Nb electrodes. A false-color scanning electron microscope (SEM) image of our device A is shown in Fig. 1a, while a device schematic is depicted in Fig. 1b along with the coordinate system. Note that the supercurrent flows across the TINW in our devices, while it flows along the wire in conventional devices [28, 29]. We observed robust supercurrents at low temperature (Fig. 1c) and the plot of differential resistance $dV/dI$ vs junction voltage $V$ (Fig. 1d) presents signatures of multiple Andreev reflection (MAR). The energy positions of these MAR peaks (Fig. 1d inset) give the SC gap energy $\Delta_{\rm MAR}\approx 0.9$ meV, which is close to the SC gap of the Nb electrodes (1.0 meV). From the $I$-$V$ characteristics, we can also extract [30, 31] the averaged junction transparency $\tau_{\rm av}$; our analysis (see supplement) gives $\tau_{\rm av}\approx$ 0.76, which is reasonably high for a TI-JJ [32]. These results demonstrate that the side-contact of a superconductor to a thin TI flake is possible and useful for the SC proximity effect. Similar side-contact technique is widely used for graphene [33, 34], but it is new to TI flakes.
Importantly, our sandwich TINW junction forms a well-defined Josephson junction along the full length. This is evidenced by the Fraunhofer patterns shown in Figs. 1e-g for three devices. Here, the magnetic field $B_{z}$ is applied perpendicular to the junction plane. The $B_{z}$-dependence of $I_{c}$ presents clear nodes, which strongly suggests that the supercurrent is not due to superconducting shorts but the SC phase difference $\varphi$ winds along the junction (in the $x$-direction) in response to $B_{z}$. In all three devices, the $B_{z}$ values at the first $I_{c}$ minimum are essentially consistent with the expected $B_{z}$ values to create the SC flux quantum $\Phi_{0}^{s}=h/2e$ in the junction area, taking into account the flux-focusing effect [32, 29, 35] (see supplement for details). We note that the observed Fraunhofer pattern is not very regular in all three devices. This is understood as a result of inhomogeneity in the critical current along the junction [32, 36, 37]. This irregularity suggests that either the side of the etched TINW is not smooth or the induced pairing potential is nonuniform, resulting in locally varying $I_{c}$.
@@ II Results
Figures 2a-c show the $I_{c}(B_{x})$ behavior thus obtained in devices A, B, and C. One can see that $I_{c}$ presents pronounced oscillations as a function of $B_{x}$, almost vanishing at the minima in devices A and C, but virtually recovers its zero-field value at the maxima. These strong oscillations of $I_{c}$ are completely different from the Fraunhofer pattern and are akin to the SQUID pattern. To understand the period $\Delta B_{x}$ of these $I_{c}$ oscillations, we evaluated the geometrical and effective cross-sectional area of a total of five TINW devices: The geometrical area $A_{\mathrm{geo}}$ is calculated with the actual width $W$ and the thickness $d$, while the effective area $A_{\rm eff}$ takes into account the penetration depth of the surface-state wavefunction; namely, we assumed that the effective electronic boundary of the nanowire is $\sim$2.5 nm deep from the outer surface [38, 39, 40].
Figure 3: Relation between the TINW cross-section and the $B_{x}$-period of the $I_{c}$ oscillations. The period $\Delta B_{x}$, defined as the separation between the first minima for positive and negative $B_{x}$, is obtained from the $I_{c}$ oscillations in devices A-E (see supplement). The geometrical cross-section $A_{\mathrm{geo}}$ is simply $W$ times $d$. The effective cross-section $A_{\rm eff}$ takes into account the depth of the surface-state wavefunction ($\sim$2.5 nm) [38, 40]. $A_{\mathrm{geo}}$ and $A_{\mathrm{eff}}$ are plotted vs $1/\Delta B_{x}$ with open and filled symbols, respectively. The solid and dashed lines are the expected relations when the period corresponds to the magnetic flux of $h/(2e)$ and $h/e$, respectively.
In Fig. 3, $A_{\mathrm{geo}}$ and $A_{\mathrm{eff}}$ are plotted vs $1/\Delta B_{x}$, along with the lines showing the area corresponding to the flux change $\Delta\Phi=h/e$ and $\Delta\Phi=h/(2e)$. One can easily see that the observed $\Delta B_{x}$ corresponds to the flux change of $h/(2e)$ if $A_{\rm eff}$ is taken. Thus, we conclude that the period of the $I_{c}$ oscillations is set by $h/(2e)$, which is expected for a SQUID. Note that the flux-focusing effect from the SC electrodes can be neglected in this geometry, because the Nb electrodes are almost fully penetrated by parallel magnetic fields. Note also that at the maximum $B_{x}$ field of this experiment, $\pm$6 T, the superconductivity of the Nb electrodes was about to be suppressed, making the $I_{c}$ lobes in Figs. 2a-c to be distorted towards 6 T.
@@
One of the advantages of our geometry is that the chemical potential $\mu$ can be tuned by a back-gate voltage $V_{\rm G}$. The $V_{\rm G}$-dependence of $I_{c}$ shown in Fig. 2d for device B demonstrates that $\mu$ can be tuned across the Dirac point, which appears to be located at around $V_{\rm G}\approx-10$ V. The $I_{c}$ increases by a factor of 3 as $\mu$ moves away from the Dirac point into the $n$-type regime, while $I_{c}$ presents a saturation in the $p$-type regime. Such a behaviour has been observed in planar JJs based on TIs [32, 43] and graphene [44], where the saturation in the $p$-type regime was attributed to the formation of a $pn$-junction between the gated junction channel and the areas that are in direct contact with the SC electrodes. The weak oscillatory $I_{c}(V_{\rm G})$ behavior in the $p$-type regime could be due to a Fabry-Perot-like interference [34, 44].
One can see in Figs. 2e-f that the pronounced $I_{c}$ oscillations as a function of $B_{x}$ are stable with respect to changes in $V_{\rm G}$, although an offset in $I_{c}$ grows with $V_{\rm G}$ in the $n$-type regime. Interestingly, device B is tuned with $V_{\rm G}\approx-20$ V to a regime where $I_{c}$ becomes zero at the minima. Such a disappearance of $I_{c}$ is possible only when there is no bulk contribution and the critical current at the top and bottom surfaces are equal. Hence, one can conclude that our TI nano-SQUID is surface-dominated.
@@
where $\varphi$ is the superconducting phase difference across the junctions and $I_{t}$ ($I_{b}$) is the maximum supercurrent through the top (bottom) surface. The critical current can be obtained by finding the maximum as a function of $\varphi$, $I_{c}=\max_{\varphi}\{I(\varphi)\}$. The resulting $I_{c}$ as a function of $\Phi$ are shown in Fig. 4a for two ratios of $I_{t}/I_{b}$. As is usual for a SQUID, $I_{c}$ exhibits large oscillations with periodicity $h/2e$, and the depth of the $I_{c}$ minima at $\Phi=h/4e$ depends on the asymmetry, $I_{t}/I_{b}$. We obtain $I_{c}\approx$ 0 at the minima in the symmetric case $I_{t}\approx I_{b}$. Therefore, the observed $V_{G}$ dependence of the $I_{c}$ oscillations shown in Figs. 2e-f is understood to be essentially a result of the gate-induced change in $I_{t}$ and $I_{b}$.
Figure 4: Results of numerical simulations: (a) The $I_{c}$ oscillations calculated for the simple phenomenological model (dotted lines) and the tight-binding model (solid line). The dominant period $\Delta\Phi=\Phi_{0}^{s}$ [$=h/(2e)$] appears in both models. In the phenomenological model Eq. (1), the ratio $I_{t}/I_{b}$ determines the size of the oscillations in $I_{c}$. The result of the tight-binding model for small $\Delta$ shows a superposition of oscillations with period $2\Phi_{0}^{s}$ [$=h/e$], which arises due to dependence of the normal state subbands on flux. This odd-even effect is not observed in our experiment, suggesting well-defined subbands is likely not formed. (b) Sketch: When the flux $\Phi$ is in the range $(n-\frac{1}{2})\Phi_{0}^{s}<\Phi<(n+\frac{1}{2})\Phi_{0}^{s}$ ($n$ odd), the columnar nano-SQUID hosts MZMs at the ends. (c) Topological phase diagram, spectral gap of top- and bottom surfaces $E_{{\rm gap,t/b}}$, and gauge-invariant equilibrium phase difference $\gamma_{0,{\rm t/b}}=\varphi_{0}-\frac{2\pi}{\Phi_{0}^{s}}\int{\bm{A}}{\rm d}{\bm{l}}=\varphi\pm\pi\Phi/\Phi_{0}^{s}$ at top- and bottom surfaces as function of magnetic flux $\Phi$ threading the nano-SQUID; calculated from Eq. (1). Topological regions are shaded in red.
To support this phenomenological nano-SQUID model, we performed numerical simulations of our set-up by using a BHZ tight-binding model of a 3D TINW with finite cross-section and of infinite length, such that the momentum $k$ along the wire is a good quantum number (see Methods). We assume a spatial dependence of the pairing potential $\Delta(y)$ induced in the TINW surface, such that there is a phase difference of $\varphi$ across the top/bottom surface; namely, we assume the SC phase of $-\varphi/2$ for the left side surface and $+\varphi/2$ for the right. The ground state value of $\varphi$ is set by the flux $\Phi$, but it changes when a transverse current flows. This model effectively describes a SQUID with two JJs at the top and bottom of the TINW surface. The original TINW states are now broken into four parts, two at the sides with a SC gap and two at the top and bottom with ABSs. By diagonalizing the Bogoliubov-de Gennes Hamiltonian for this tight-binding model, we calculate the total energy $E$ as a function of $\varphi$, and its derivative gives the CPR, which allows us to calculate $I_{c}$ (see Methods and supplement).

### v3

# Topological Insulator nano-SQUID: Flux-tunable platform for topological superconductivity

@@ Abstract
Many efforts have been made in the past decade to realize topological superconductivity using superconducting proximity effect, but an ideal platform is still lacking. A 3D topological insulator (TI) is promising for this purpose due to the spin-momentum-locked surface state. Here we propose a novel yet simple TI platform which gives rise to a topological phase that is robust against disorder. It consists of a bulk-insulating rectangular TI nanowire laterally sandwiched by two superconductors. In this structure, the top and bottom surfaces individually work as SNS line junctions, forming a nanometer-scale columnar SQUID in which the nanowire cross-section defines the threading magnetic flux $\Phi$ in axial magnetic fields. We theoretically show that, when the two junctions are asymmetric, a robust topological phase occurs periodically for a wide range of $\Phi$, independently of the chemical potential. Our experiment found that a TI device of this structure indeed behaves as a columnar nano-SQUID where the supercurrent flows only through the top and bottom surfaces with vanishing bulk contribution. Furthermore, the top/bottom asymmetry can be tuned by a back gate, a key ingredient for the topological phase.
@@ I Introduction
The main feature of three-dimensional (3D) topological insulators (TIs) is the existence of gapless surface Dirac fermions that are protected by time-reversal symmetry [1, 2, 3]. These Dirac fermions are spin-non-degenerate with the spin direction dictated by the momentum; this spin-momentum locking is the key ingredient for realizing “spinless” topological superconductivity that gives rise to Majorana zero-modes (MZMs) [4]. The non-Abelian statistics obeyed by MZMs allows for topological quantum computing based on braiding operations [5], so TIs have a high potential for such applications. However, TIs have been left behind in the main-stream efforts to realize topological superconductivity, largely due to the less developed technologies for TI device fabrications compared to semiconductor devices [6]. Unfortunately, despite significant efforts in the main-stream semiconductor platforms, the generation of MZMs is still elusive [7, 8]. This is mainly due to the sensitivity of the semiconductor platforms to Coulomb disorder [9, 8, 10, 11], and a new platform that is inherently insensitive to chemical-potential fluctuations is strongly called for.
@@
Such a topological 1D Majorana state should in principle have a MZM when it has an end, but if the line junction ends in the middle of the TI surface, the MZM will hybridize with the gapless Dirac fermions in the rest of the surface; therefore, Fu and Kane proposed a phase-biased tri-junction to generate a MZM that is protected by a gap [4]. A Josephson vortex in a TI line junction also hosts a MZM due to the naturally-occurring phase bias [12, 13]. These MZMs generated in a TI line junction have the important advantage that they are disorder-resilient [9], as their existence does not depend on the chemical potential at all, provided that the TI is bulk-insulating.
@@
Figure 1: Sandwich TINW Josephson junction. (a) False-color SEM image of device A; an exfoliated TI flake is dry-etched to leave a NW (pink), and the etched area is subsequently filled with Nb (green), such that the TINW is side-contacted by superconducting Nb. There are remaining TI pieces that do not play any role. (b) Schematic of our device: Back-gated TINW is laterally sandwiched between two superconducting Nb electrodes to a Josephson junction. In- and out-of-plane magnetic fields can be applied. (c) $dV/dI$ vs $I_{\rm b}$ data in device A at various $T$ shown for zero back-gate voltage; inset shows the $T$-dependence of the critical current $I_{c}$ extracted from these data. We define $I_{c}$ as the bias current above which $dI/dV$ exceeded 50 $\Omega$. (d) Differential resistance $dV/dI$ vs junction voltage $V$ for device A in $B$ = 0 T. Red arrows mark the position of the peaks due to multiple Andreev reflection (MAR). We assign the index $n$ from 1 to 3 to the peaks ($n$ = 1 for the largest $V$). Inset shows a plot of the $V$ value for index $n$ (which we call $V_{n}$) vs $1/n$. (e-g) Color mappings of $dV/dI$ as a function of $I_{\mathrm{b}}$ and $B_{\mathrm{z}}$ for devices A, B and C measured without any in-plane magnetic field at $T=30$ mK, presenting Fraunhofer patterns. The critical current manifests itself as a bright orange line to border the black area corresponding to zero resistance. Irregular Fraunhofer features are attributed to an inhomogeneous supercurrent density along the junction as demonstrated in supplementary Fig. S5
@@
To exploit the benefits of the TI line junction in an alternative structure, we fabricate unconventional devices in which a rectangular nanowire, etch-fabricated from a bulk-insulating TI flake [14], is side-contacted by a superconductor on both left and right sides, creating a lateral sandwich SNS junction with a TI nanowire (TINW) as the N part (see Fig. 1b). In such a sandwich junction, if the TINW is bulk-insulating, the supercurrent flows only through the top and bottom surfaces (and a very small portion through the ends of the TINW) via the ABSs formed there. Since the axial magnetic field applied parallel to the TINW creates a magnetic flux $\Phi$ which controls the phase difference $\varphi$ at the top and bottom JJs, this structure can be viewed as a nanometer-scale columnar SQUID (superconducting quantum interference device) where the loop area corresponds to the TINW cross-section. A similar nano-SQUID was previously realized with a double-layer graphene heterostructure [15]. Our device has a bottom gate, which allows us to control the chemical potential in the TINW as well as to introduce asymmetry between the top and bottom junctions. A similar lateral sandwich junction was previously realized with a TI by using a diffusion process [16], but the device was not gate-tunable and it was not clear if the supercurrent flows only through the surface or also through the bulk.
@@
In our devices, we find that the critical current $I_{c}$ of the device exhibits pronounced oscillations as a function of $\Phi$ with a period $h/(2e)$ and, in certain situations, $I_{c}$ is almost completely suppressed when $\Phi=h/(4e)$ but recovers almost entirely for $\Phi=h/(2e)$. This is exactly the behavior expected for a nano-SQUID with completely surface-dominated supercurrent with negligible bulk contribution. Theoretically, we first discuss a phenomenological SQUID model to explain the observed behavior, and then present numerical simulations based on a microscopic model of two line junctions made on a TINW surface, which support the SQUID model. Furthermore, we theoretically show that when the SQUID is asymmetric, the magnetic flux $\Phi$ causes only one of the two line junctions to go through the topological phase transition of the Fu-Kane type [4] and the SQUID is topological for $(n-\frac{1}{2})\Phi_{0}^{s}<\Phi<(n+\frac{1}{2})\Phi_{0}^{s}$ with odd-integer $n$. This topological phase, which occurs for a wide parameter range, is robust against chemical-potential fluctuations due to disorder. The columnar nano-SQUID structure has an advantage over tri-junction or Josephson-vortex structures in that the MZMs appear at the ends of the TINW and are easier to access. Note that there have also been related theoretical proposals to use the quasi-one-dimensional states of TINWs under superconducting proximity effect for generating MZMs [17, 18, 19, 20, 21]; however, these rely on the realization of a well-defined subband structure, i.e., that electron states on the top and bottom surfaces are coherently connected, which is most likely not realized in our device.
@@
II Columnar nano-SQUID
Device characterization: Our TINWs are based on bulk-insulating BiSbTeSe2 [22, 23] and are side-contacted by Nb electrodes. A false-color scanning electron microscope (SEM) image of our device A is shown in Fig. 1a, while a device schematic is depicted in Fig. 1b along with the coordinate system. Note that the supercurrent flows across the TINW in our devices, while it flows along the wire in conventional devices [24, 25]. We observed robust supercurrents at low temperature (Fig. 1c) and the plot of differential resistance $dV/dI$ vs junction voltage $V$ (Fig. 1d) presents signatures of multiple Andreev reflection (MAR). The energy positions of these MAR peaks (Fig. 1d inset) give the SC gap energy $\Delta_{\rm MAR}\approx 0.9$ meV, which is close to the SC gap of the Nb electrodes (1.0 meV). From the II-$V$ characteristics, we can also extract [26, 27] the averaged junction transparency $\tau_{\rm av}$; our analysis (see supplement) gives $\tau_{\rm av}\approx$ 0.76, which is reasonably high for a TI-JJ [28]. These results demonstrate that the side-contact of a superconductor to a thin TI flake is possible and useful for the SC proximity effect. Similar side-contact technique is widely used for graphene [29, 30], but it is new to TI flakes.
Importantly, our sandwich TINW junction forms a well-defined Josephson junction along the full length. This is evidenced by the Fraunhofer patterns shown in Figs. 1e-g for three devices. Here, the magnetic field $B_{z}$ is applied perpendicular to the junction plane. The $B_{z}$-dependence of $I_{c}$ presents clear nodes, which strongly suggests that the supercurrent is not due to superconducting shorts but the SC phase difference $\varphi$ winds along the junction (in the $x$-direction) in response to $B_{z}$. In all three devices, the $B_{z}$ values at the first $I_{c}$ minimum are essentially consistent with the expected $B_{z}$ values to create the SC flux quantum $\Phi_{0}^{s}=h/2e$ in the junction area, taking into account the flux-focusing effect [28, 25, 31] (see supplement for details). We note that the observed Fraunhofer pattern is not very regular in all three devices. This is understood as a result of inhomogeneity in the critical current along the junction [28, 32, 33]. This irregularity suggests that either the side of the etched TINW is not smooth or the induced pairing potential is nonuniform, resulting in locally varying $I_{c}$. In the supplement, we show that the observed irregular Fraunhofer pattern can actually be reproduced with a position-dependent $I_{c}$ determined by a maximum-entropy fitting [28]. The weak asymmetry between $\pm B_{z}$ is most likely an additional self-field effect [34].
@@ II Columnar nano-SQUID
Figures 2a-c show the $I_{c}(B_{x})$ behavior thus obtained in devices A, B, and C, that are slightly different in dimensions; device C was thicker and wider than others. One can see that $I_{c}$ presents pronounced oscillations as a function of $B_{x}$, almost vanishing at the minima in devices A and C, but virtually recovers its zero-field value at the maxima. These strong oscillations of $I_{c}$ are completely different from the Fraunhofer pattern and are akin to the SQUID pattern. To understand the period $\Delta B_{x}$ of these $I_{c}$ oscillations, we evaluated the geometrical and effective cross-sectional area of a total of five TINW devices: The geometrical area $A_{\mathrm{geo}}$ is calculated with the actual width $W$ and the thickness $d$, while the effective area $A_{\rm eff}$ takes into account the penetration depth of the surface-state wavefunction; namely, we assumed that the effective electronic boundary of the nanowire is $\sim$2.5 nm deep from the outer surface [35, 36, 37].
Figure 3: Relation between the TINW cross-section and the $B_{x}$-period of the $I_{c}$ oscillations. The period $\Delta B_{x}$, defined as the separation between the first minima for positive and negative $B_{x}$, is obtained from the $I_{c}$ oscillations in devices A-E (see supplement). The geometrical cross-section $A_{\mathrm{geo}}$ is simply $W$ times $d$. The effective cross-section $A_{\rm eff}$ takes into account the depth of the surface-state wavefunction ($\sim$2.5 nm) [35, 37]. $A_{\mathrm{geo}}$ and $A_{\mathrm{eff}}$ are plotted vs $1/\Delta B_{x}$ with open and filled symbols, respectively. Error bars represent the maximum estimated uncertainty; assuming uncertainties of 1 nm in $d$, 5 nm in $W$, and 100-200 mT (depending on the device) in the magnetic field for $I_{c}$ minima.
The solid and dashed lines are the expected relations when the period corresponds to the magnetic flux of $h/(2e)$ and $h/e$, respectively.
In Fig. 3, $A_{\mathrm{geo}}$ and $A_{\mathrm{eff}}$ are plotted vs $1/\Delta B_{x}$, along with the lines showing the area corresponding to the flux change $\Delta\Phi=h/e$ and $\Delta\Phi=h/(2e)$. One can easily see that the observed $\Delta B_{x}$ corresponds to the flux change of $h/(2e)$ if $A_{\rm eff}$ is taken. Thus, we conclude that the period of the $I_{c}$ oscillations is set by $h/(2e)$, which is expected for a SQUID. Note that the flux-focusing effect from the SC electrodes can be neglected in this geometry, because the Nb electrodes are almost fully penetrated by parallel magnetic fields. Note also that at the maximum $B_{x}$ field of this experiment, $\pm$6 T, the superconductivity of the Nb electrodes was about to disappear, making the $I_{c}$ lobes in Figs. 2a-c to be distorted and suppressed towards 6 T. The reasonable agreement of $A_{\rm eff}$ with the data suggests that possible fabrication damages to the surface are benign enough not to affect the surface-state penetration depth.
@@
One of the advantages of our geometry is that the chemical potential $\mu$ can be tuned by a back-gate voltage $V_{\rm G}$. The $V_{\rm G}$-dependence of $I_{c}$ shown in Fig. 2d for device B demonstrates that $\mu$ can be tuned across the Dirac point, which appears to be located at around $V_{\rm G}\approx-10$ V. The $I_{c}$ increases by a factor of 3 as $\mu$ moves away from the Dirac point into the $n$-type regime, while $I_{c}$ presents a saturation in the $p$-type regime. Such a behaviour is characteristic of gapless Dirac fermions and has been observed in planar JJs based on TIs [28, 40] and graphene [41]; the saturation in the $p$-type regime was attributed to the formation of a $pn$-junction between the gated junction channel and the areas that are in direct contact with the SC electrodes. The weak oscillatory $I_{c}(V_{\rm G})$ behavior in the $p$-type regime could be due to a Fabry-Perot-like interference [30, 41].
One can see in Figs. 2e-f that the pronounced $I_{c}$ oscillations as a function of $B_{x}$ are stable with respect to changes in $V_{\rm G}$, although an offset in $I_{c}$ grows with $V_{\rm G}$ in the $n$-type regime. Interestingly, device B is tuned with $V_{\rm G}\approx-20$ V to a regime where $I_{c}$ becomes zero at the minima. Such a disappearance of $I_{c}$ is possible only when there is no bulk contribution and the critical current at the top and bottom surfaces are equal. Hence, one can conclude that our TI nano-SQUID is surface-dominated. This “surface only” supercurent combined with its $V_{\rm G}$ dependence that is characteristic of gapless Dirac fermions are a clear manifestation of the TI nature. Note that the formation of charge puddles in the bulk due to disorder is efficiently suppressed by the shielding effect of the superconducting electrodes [42], which is an additional advantage of the nano-SQUID.
Figure 4: Results of numerical simulations: (a) The $I_{c}$ oscillations calculated for the simple phenomenological model (dotted lines) and the tight-binding model (solid line). The dominant period $\Delta\Phi=\Phi_{0}^{s}$ [=h/(2e)=h/(2e)] appears in both models. In the phenomenological model Eq. (1), the ratio $I_{t}/I_{b}$ determines the size of the oscillations in $I_{c}$. The result of the tight-binding model for small $\Delta$ shows a superposition of oscillations with period $2\Phi_{0}^{s}$ [=h/e=h/e], which arises due to dependence of the normal state subbands on flux. This odd-even effect is not observed in our experiment, suggesting well-defined subbands is likely not formed. (b) Sketch: When the flux $\Phi$ is in the range $(n-\frac{1}{2})\Phi_{0}^{s}<\Phi<(n+\frac{1}{2})\Phi_{0}^{s}$ ($n$ odd), the columnar nano-SQUID hosts MZMs at the ends. (c) Topological phase diagram, spectral gap of top- and bottom surfaces $E_{{\rm gap,t/b}}$, and gauge-invariant equilibrium phase difference $\gamma_{0,{\rm t/b}}=\varphi_{0}-\frac{2\pi}{\Phi_{0}^{s}}\int{\bm{A}}{\rm d}{\bm{l}}=\varphi\pm\pi\Phi/\Phi_{0}^{s}$ at top- and bottom surfaces as function of magnetic flux $\Phi$ threading the nano-SQUID; calculated from Eq. (1). Topological regions are shaded in red.
@@
where $\varphi$ is the superconducting phase difference across the junctions and $I_{t}$ ($I_{b}$) is the maximum supercurrent through the top (bottom) surface. The critical current can be obtained by finding the maximum as a function of $\varphi$, $I_{c}=\max_{\varphi}\{I(\varphi)\}$. The resulting $I_{c}$ as a function of $\Phi$ are shown in Fig. 4a for two ratios of $I_{t}/I_{b}$. As is usual for a SQUID, $I_{c}$ exhibits large oscillations with periodicity $h/2e$, and the depth of the $I_{c}$ minima at $\Phi=h/4e$ depends on the asymmetry, $I_{t}/I_{b}$. We obtain $I_{c}\approx$ 0 at the minima in the symmetric case $I_{t}\approx I_{b}$. Therefore, the observed $V_{G}$ dependence of the $I_{c}$ oscillations shown in Figs. 2e-f is understood to be essentially a result of the gate-induced change in $I_{t}$ and $I_{b}$. For example, the data in Fig. 2f for $V_{\rm G}$ = 20 V and 20-20 V correspond to $I_{t}/I_{b}\simeq$ 0.5 and 1, respectively.
To support this phenomenological nano-SQUID model, we performed numerical simulations of our set-up by using a BHZ tight-binding model of a 3D TINW with finite cross-section and of infinite length, such that the momentum $k$ along the wire is a good quantum number (see Methods). We assume a spatial dependence of the pairing potential $\Delta(y)$ induced in the TINW surface, such that there is a phase difference of $\varphi$ across the top/bottom surface; namely, we assume the SC phase of $\varphi/2$ for the left side surface and $\varphi/2$ for the right. The ground state value of $\varphi$ is set by the flux $\Phi$, but it changes when a transverse current flows. This model effectively describes a SQUID with two JJs at the top and bottom of the TINW surface. The original TINW states are now broken into four parts, two at the sides with a SC gap and two at the top and bottom with ABSs. By diagonalizing the Bogoliubov-de Gennes Hamiltonian for this tight-binding model, we calculate the total energy $E$ as a function of $\varphi$, and its derivative gives the CPR, which allows us to calculate $I_{c}$ (see Methods and supplement).